# BayeSN and SALT: A Comparison of Dust Inference Across SN Ia Light-curve Models with DES5YR

### v2

@@ Abstract
In recent years there has been significant debate around the impact of dust on SNe Ia, a major source of uncertainty in cosmological analyses. We perform the first cross-comparison of the probabilistic hierarchical SN Ia SED model BayeSN with the conventional SALT model, an important test given the history of conflicting conclusions regarding the distributions of host galaxy dust properties between the two. Applying BayeSN to SALT-based simulations, we find that BayeSN is able to accurately recover our simulated inputs, establishing excellent consistency between the two models. When inferring dust parameters with simulated samples including non-Ia contamination, we find that our choice of photometric classifier causes a bias in the inferred dust distribution; this arises because SNe Ia heavily impacted by dust are misclassified as contaminants and excluded. We then apply BayeSN to the sample of SNe from DES5YR to jointly infer host galaxy dust distributions and intrinsic differences on either side of the ‘mass step’ at $10^{10}$ Mdirect-$\odot$. We find evidence in favour of an intrinsic contribution to the mass step and differing $R_{V}$ distributions. We also build on recent results supporting an environmental-dependence on the secondary maximum of SNe Ia in $i$-band. Twenty days post-peak, we find an offset in intrinsic $i$-band light curve between each mass bin at a significance in excess of $3\sigma$.
@@ 1 Introduction
Type Ia supernovae (SNe Ia) were crucial tools in discovering the accelerating expansion of the universe (Riess et al., 1998; Perlmutter et al., 1999). A possible cause of this expansion, dubbed ‘dark energy’, has been further constrained in subsequent years; Popovic et al. (2023a); Vincenzi et al. (2024); Brout et al. (2022) have measured the dark energy equation-of-state $w$ with increasing precision. While the current constraints on $w$ are statistically-dominated, future surveys such as the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b), the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) and Roman Space Telescope promise orders of magnitude more supernovae, placing a greater need on more precise control of systematic uncertainties.
@@
To estimate N Ia distances with the precision needed to constrain cosmological parameters, a standardisation process is required. Without standardisation, SN Ia peak-optical absolute magnitudes have a natural scatter of ${\sim}0.40$ mag; however, two empirical relationships have been found that reduce this scatter to ${\sim}0.15$ mag. The first is the Phillips relation, from Phillips (1993); Tripp (1998), that shows that brighter supernovae have a slower decline rate post-peak; this relation, combined with the discovery that bluer SNe Ia are intrinsically brighter (Riess et al., 1996), are crucial to the usage of SNe Ia in the modern day. Most modern cosmological analyses use the SALT light-curve fitter Guy et al. (2010); Kenworthy et al. (2021).
@@
However, there are some shortcomings to the SALT approach. The process of standardisation within SALT is done through the use of the Tripp formula, a linear relation of magnitude vs. light curve stretch $x_{1}$ (i.e. the Phillips relation) and a colour parameter $c$ that roughly corresponds to peak $B-V$ apparent colour. SALT assumes a single colour law for all sources of SN Ia reddening and does not disentangle intrinsic colour variation from dust reddening. These two separate effects are therefore captured by a single linear coefficient for the entire population of SNe Ia; analyses of the host galaxy dust properties of SNe Ia have demonstrated that this approximation is not physical (e.g. Mandel et al., 2017; Brout & Scolnic, 2021; Popovic et al., 2023b; Thorp et al., 2021; Ward et al., 2023; Grayling et al., 2024).
The current largest systematic uncertainties impacting cosmology with SNe Ia are astrophysical (Popovic et al., 2023a; Vincenzi et al., 2024), and decreasing these uncertainties will require further understanding of the complex astrophysical processes that drive SNe Ia. Dhawan et al. (2024) recently explored how a variety of effects, including evolution of SN Ia host galaxy dust properties with redshift, would lead to offsets from a fiducial $\Lambda$CDM cosmology. In the past decade, new discoveries have improved our understanding of how SNe Ia relate to their host galaxies. The ‘mass step’, the observation that SNe Ia in higher mass galaxies are brighter than their low mass counterparts post-standardisation, was discovered by Sullivan et al. (2010); Kelly et al. (2010). Since the initial discovery of the mass step, it has been consistently observed in a variety of independent optical samples (Kelly et al., 2010; Sullivan et al., 2010; Lampeitl et al., 2010; Childress et al., 2014; Betoule et al., 2014; Jones et al., 2018; Smith et al., 2020; Kelsey et al., 2021a). This observation has been further extended to other host-galaxy environmental properties, including star formation rate (SFR), specific star formation rate (sSFR), and host-galaxy colour (Rigault et al., 2018; Briday et al., 2022; Kelsey et al., 2023). These differences are often stronger when considering properties of the local environments of SNe Ia rather properties derived from the entire host galaxy (e.g. Rigault et al., 2013, 2015, 2020; Jones et al., 2015; Moreno-Raya et al., 2016a, b; Jones et al., 2018; Kim et al., 2018; Roman et al., 2018; Kim et al., 2019; Rose et al., 2019; Kelsey et al., 2021b).
From the SALT side, Brout & Scolnic (2021) and later Popovic et al. (2023b); Wiseman et al. (2022) have had success in attributing the root cause of the mass step to dust. These papers showed that an intrinsic distribution of SN Ia colours, reddened and dimmed by dust from the host-galaxy, is able to recreate observed features of SN Ia data. Specifically, these works postulated that the mass step is a result of a difference in the distribution of SN Ia line-of-sight $R_{V}$ values between high- and low-mass galaxies. This dust-based scatter model was used in the Brout et al. (2022); Vincenzi et al. (2024); Popovic et al. (2023a) cosmological analyses.
@@
Given the differing conclusions regarding dust from SALT and BayeSN analyses, it is important to understand exactly where these differences arise and try to reconcile these results. In this work, we perform the first cross-comparison between the two by applying BayeSN for inference of dust properties on SALT-based simulations similar to those used in the DES5YR cosmology analysis. Following from this, we apply BayeSN to a sample of SNe Ia from the DES5YR sample to comment on the relative contributions of dust and intrinsic effects in explaining the mass step. We also compare these results on the real DES5YR sample to those from SALT-based analyses.
@@ 2.1 BayeSN
BayeSN uses the optical-NIR SN Ia SED template from Hsiao et al. (2007) as the base SED, defined as $S_{0}(t,\lambda_{r})$, with an arbitrary scaling factor $M_{0}$ of $-19.5$. The rest-frame wavelength and rest-frame phase relative to maximum flux are given by $\lambda_{r}$ and $t$ respectively. Latent variables which are different for each SN are denoted with s; the other parameters are global hyperparameters shared across the whole population. The different components which make up the model are as follows:
@@
  • $\theta_{1}^{s}$ is a principal component coefficient which quantifies the effect of the $W_{1}$ FPC for each supernova. In concert with $W_{1}(t,\lambda_{r})$, this accounts for the ‘broader-brighter’ relationship observed across SNe Ia Phillips (1993). $\theta_{1}$ is defined to follow a normal distribution such that $\theta_{1}\sim\mathcal{N}(0,1)$.
@@ 2.1.1 BayeSN Treatment of the Mass Step
$R_{V}^{s}\sim\begin{cases}TN(\mu_{R_{V},\text{HM}},\sigma_{R,\text{HM}}^{2},1.2,\infty),&\text{ if }M_{*}^{s}>M_{\text{split}}\\ TN(\mu_{R_{V},\text{LM}},\sigma_{R,\text{LM}}^{2},1.2,\infty),&\text{ if }M_{*}^{s}<M_{\text{split}}\\ \end{cases}$ (2)
@@
$\delta M^{s}\sim\begin{cases}N(0,\sigma_{0,\text{HM}}^{2}),&\text{ if }M_{*}^{s}>M_{\text{split}}\\ N(\Delta M_{0},\sigma_{0,\text{LM}}^{2}),&\text{ if }M_{*}^{s}<M_{\text{split}}.\\ \end{cases}$ (3)
@@ 2.1.2 Conditioning on Distance
For this work, in our analysis of the DES5YR sample we assume a flat $\Lambda$CDM cosmology with $\Omega_{m}=0.28$ and $H_{0}$ $=$ 73.24km s-1 Mpc-1 for consistency with previous BayeSN analyses. For our inference on SALT-based simulations, however, we assume a flat $\Lambda$CDM cosmology with $\Omega_{m}=0.311$ and $H_{0}=70$ km s-1 in order to match the cosmology assumed when creating the simulations.
@@ 2.2 Simulations with SALT
Here, we focus on the compatibility of BayeSN and SALT. To-date, there has been no cross-comparison between BayeSN and SALT-based approaches to dust inference. Given the history of conflicting results between the dust parameters derived from BayeSN (Mandel et al., 2022; Thorp et al., 2021; Grayling et al., 2024) and those from SALT (Brout & Scolnic, 2021; Wiseman et al., 2022; Popovic et al., 2023b), in particular Dust2Dust Popovic et al. (2023b), we focus on the ability of BayeSN to recover dust parameters from SALT-based simulations of SN Ia light-curves, and additionally, contaminated samples. The self-consistency of Dust2Dust is reviewed in Popovic et al. (2023b). A companion paper, Chen et al. (in prep.), focuses on a direct comparison between SALT and BayeSN distances and supernova standardisation parameters.
@@
We aim to perform a consistency check between SALT and BayeSN modeling of dust parameters, and therefore our base source model is the SALT model (Guy et al., 2007), updated by Kenworthy et al. (2021). Compared to BayeSN, SALT models the flux as
@@
which results in a similar SED-model: $M_{0}$ and $M_{1}$ roughly correspond to $W_{0}$ and $W_{1}$ in BayeSN, modelling the average SED and the first-order mode of variation to the SED respectively. The colour law $CL(\lambda)$ captures the colour-related variation in the SED across the population, combining intrinsic variation with the effect of host galaxy dust extinction which are respectively modelled by $\mathbf{\Sigma_{\epsilon}}$ and $\xi$ in BayeSN. SALT $x_{1}^{s}$ finds its BayeSN analogue in $\theta^{s}_{1}$, while $c$ is roughly comparable though not analogous to $A^{s}_{V}$. In SALT, the overall amplitude of the light-curve at peak brightness, $x_{0}^{s}$, has no direct correspondent in BayeSN, which infers the distance jointly with the other parameters. $x_{1}^{s}$, $c^{s}$, and $x_{0}^{s}$ are all determined for each supernova, whereas $M_{0}$ and $M_{1}$ are population-level SEDs that are determined during the SALT training process.
@@
The dust-modeling framework introduced in Mandel et al. (2017); Brout & Scolnic (2021); Popovic et al. (2023b) makes some changes to the SALT and Tripp formulations. The distribution of SN Ia colours is comprised of a combination of a intrinsic, dust-free distribution - $c_{\rm int}$ and a reddening component due to dust ($E_{\rm dust}$). The $c$ component in Equation 5 is then
@@
2.3 Differences between BayeSN and Dust2Dust
@@
  • BayeSN assumes a constant, achromatic scatter $\delta M^{s}$ around the distance-redshift relation which follows a Gaussian distribution, in addition to a chromatic scatter term. SNANA simulations assume no specific distribution in this achromatic scatter when simulating, though it is sufficiently described by a Gaussian.
@@ 3 Data
We make use of the Dark Energy Survey ‘5-year’ data release (DES-SN, (Flaugher et al., 2015; Sánchez et al., 2024). The DES observing program ran for 5 years, observing likely SNe Ia from $1.130.1<z<$ 1.13 with $griz$ bands. The host galaxy information for the DES supernovae is taken from co-added images (Wiseman et al., 2020), with environmental properties such as stellar mass and colour fit from the spectral energy distributions by Smith et al. (2020); Kelsey et al. (2023). A spectroscopic follow-up program by the Australian DES survey, OzDES (Childress et al., 2017; Lidman et al., 2020), provided spectroscopic redshifts of the host galaxies, which was paired with the state-of-the-art photometric classification program SuperNNova from Möller & de Boissière (2020). The resulting DES-SN sample is the largest single-telescope sample to-date, comprised of 1500 likely SNe Ia. A more comprehensive review of the SNe Ia used in the DES 5-year analysis can be found in Vincenzi et al. (2024); Möller et al. (2024).
@@ 3.1 Establishing the BayeDESN Sample
We additionally consider an alternative sample, the ‘Cosmo’ sample, which follows these requirements but additionally includes only SNe in the final DES5YR cosmology sample. The specific cuts applied to these samples, and the number of SNe they respectively cut, are detailed in Table 1.
@@ 4 Methodology
The focus of this work is to compare BayeSN and SALT, with specific interest to the compatibility of their approaches to dust inference. To do so, we perform a suite of tests with simulated data, as detailed in sub-sections 4.1 and 4.2. Then, in Section 6, wepply BayeSN to the real BayeDESN sample, to infer differences in dust and intrinsic properties on either side of the mass step.
@@ 4.1 Recovery of Simulated Parameters
Following Section 2.2, we simulate 50 simulacra of the DES-5YR dataset, varying supernova and dust properties according to Table 2. These simulations are then fit using the BayeSN model for hierarchical inference. We infer all global parameters including $W_{0},W_{1},\Sigma_{\epsilon}$ as well as the mean and standard deviation of the $R_{V}$ population, the population mean extinction $\tau_{A}$, and an achromatic intrinsic scatter $\sigma_{0}$; this allows the BayeSN model the flexibility to capture the variation in the SALT SED model while simultaneously inferring dust properties. In the cases where our simulated dust properties vary with host galaxy mass (see Table 2), we similarly split our hyperparameters to match the simulated ‘step’ at $10^{10}M_{*}$, where our dust distributions change. In these cases, we apply the ‘intrinsic SED difference’ model discussed in Section 2.1.1.
Systematic Test $R_{V}$ $E(B-V)$ $\beta_{\rm SN}$
Single $R_{V}$, $E(B-V)$ $\mathcal{N}(3,0.5)$ $\exp(0.080)$ 1
Double $R_{V}$, Single $E(B-V)$ $\mathcal{N}(2,0.5),\mathcal{N}(3,0.5)$ $\exp(0.080)$ 1
Double $R_{V}$, $E(B-V)$ $\mathcal{N}(2,0.5),\mathcal{N}(3,0.5)$ $\exp(0.125)$, $\exp(0.080)$ 1

### v3

@@ Abstract
[…]
@@ 1 Introduction
Type Ia supernovae (SNe Ia) were crucial tools in discovering the accelerating expansion of the universe (Riess et al., 1998; Perlmutter et al., 1999). A possible cause of this expansion, dubbed ‘dark energy’, has been further constrained in subsequent years; Popovic et al. (2023a); Vincenzi et al. (2024); Brout et al. (2022) have measured the dark energy equation-of-state $w$ with increasing precision. While the current constraints on $w$ are statistically-dominated, current and future surveys such as the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b), the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) and Roman Space Telescope (Roman; Spergel et al., 2015; Rose et al., 2025) promise orders of magnitude more supernovae, placing a greater need on more precise control of systematic uncertainties.
@@
To estimate SN Ia distances with the precision needed to constrain cosmological parameters, a standardisation process is required. Without standardisation, SN Ia peak-optical absolute magnitudes have a natural scatter of ${\sim}0.40$ mag; however, two empirical relationships have been found that reduce this scatter to ${\sim}0.15$ mag. The first is the Phillips relation, from Phillips (1993); Tripp (1998), that shows that brighter supernovae have a slower decline rate post-peak; this relation, combined with the discovery that bluer SNe Ia are intrinsically brighter (Riess et al., 1996), are crucial to the usage of SNe Ia in the modern day. Most modern cosmological analyses estimate distances using the SALT light-curve fitter Guy et al. (2007, 2010); Kenworthy et al. (2021).
@@
However, there are some shortcomings to the SALT approach. The process of standardisation within SALT is done through the use of the Tripp formula, a linear relation of magnitude vs. light curve stretch $x_{1}$ (i.e. the Phillips relation) and vs. colour parameter $c$ that roughly corresponds to peak BVB-V apparent colour. SALT assumes a single colour law for all sources of SN Ia reddening and does not disentangle intrinsic colour variation from dust reddening. These two separate effects are therefore captured by a single linear coefficient for the entire population of SNe Ia; analyses of the host galaxy dust properties of SNe Ia have demonstrated that this approximation is not physical (e.g. Mandel et al., 2017; Brout & Scolnic, 2021; Popovic et al., 2023b; Ward et al., 2023; Grayling et al., 2024; Duarte et al., 2023; Meldorf et al., 2023; Hallgren et al., 2025).
@@
From the SALT side, Brout & Scolnic (2021) and later Popovic et al. (2023b) have had success in attributing the root cause of the mass step to dust. These papers showed that a single population of intrinsic SN Ia colours, reddened and dimmed by dust from the host-galaxy, is able to recreate observed features of SN Ia data. Specifically, these works postulated that the mass step is a result of a difference in the distribution of SN Ia line-of-sight $R_{V}$ values between high- and low-mass galaxies, with a difference in the mean $R_{V}$ value $\mu_{R_{V}}$ in each environment of $\Delta\mu_{R_{V}}\sim 1.2-1.6$.. This dust-based scatter model was used in the Brout et al. (2022); Vincenzi et al. (2024); Popovic et al. (2023a) cosmological analyses.
[…]
Given the differing conclusions regarding dust – specifically, differences in $R_{V}$ distributions in different environments – from SALT and BayeSN analyses, it is important to understand exactly where these differences arise and try to reconcile these results. In this work, we perform the first comparison between the two by applying BayeSN for inference of dust properties on SALT-based simulations similar to those used in the DES5YR cosmology analysis. Following from this, we apply BayeSN to a sample of SNe Ia from the DES5YR sample to comment on the relative contributions of dust and intrinsic effects in explaining the mass step. We also compare these results on the real DES5YR sample to those from SALT-based analyses.
@@ 2.1 BayeSN
BayeSN uses the optical-NIR SN Ia SED template from Hsiao et al. (2007) as the base SED, defined as $S_{0}(t,\lambda_{r})$, with an arbitrary scaling factor $M_{0}$ of 19.5-19.5. The rest-frame wavelength and rest-frame phase relative to maximum flux are given by $\lambda_{r}$ and $t$ respectively. Latent variables which are different for each SN are denoted with s; the other parameters are global hyperparameters shared across the whole population. The different components which make up the model are as follows:
@@
  • $\theta_{1}^{s}$ is a principal component coefficient which quantifies the effect of the $W_{1}$ functional principal component (FPC) for each supernova. In concert with $W_{1}(t,\lambda_{r})$, this accounts for the ‘broader-brighter’ relationship observed across SNe Ia Phillips (1993). $\theta_{1}$ is defined to follow a normal distribution such that $\theta_{1}\sim\mathcal{N}(0,1)$.
@@ 2.1.1 BayeSN Treatment of the Mass Step
[…]
@@
[…]
@@ 2.1.2 Conditioning on Distance
[…]
@@ 2.2 Simulations with SALT
Here, we focus on the compatibility of BayeSN and SALT. To-date, there has been no cross-comparison between BayeSN and SALT-based approaches to dust inference. Given the history of conflicting results between the dust parameters derived from BayeSN (Mandel et al., 2022; Thorp et al., 2021; Grayling et al., 2024) and those from SALT (Brout & Scolnic, 2021; Wiseman et al., 2022; Popovic et al., 2023b), in particular Dust2Dust (Popovic et al., 2023b), we focus on the ability of BayeSN to recover dust parameters from SALT-based simulations of SN Ia light-curves, and additionally, contaminated samples. The self-consistency of Dust2Dust is reviewed in Popovic et al. (2023b).
@@
We aim to validate the perfomance of BayeSN and examine differences between the SALT and BayeSN treatments of dust, and therefore our base source model is the SALT model (Guy et al., 2007), updated to SALT3 by Kenworthy et al. (2021). Compared to BayeSN, SALT models the flux as
@@
which results in a similar SED-model: $M_{0}$ [^2] BayeSN uses the Hsiao 07 template as its base for $W_{0}$, SALT does not use a pre-existing template as the basis for $M_{0}$. and $M_{1}$ roughly correspond to $W_{0}$ and $W_{1}$ in BayeSN, modelling the average SED and the first-order mode of variation to the SED respectively. The colour law $CL(\lambda)$ captures the colour-related variation in the SED across the population, combining intrinsic variation with the effect of host galaxy dust extinction which are respectively modelled by $\mathbf{\Sigma_{\epsilon}}$ and $\xi$ in BayeSN. SALT $x_{1}^{s}$ finds its BayeSN analogue in $\theta^{s}_{1}$, while $c$ is roughly comparable though not analogous to $A^{s}_{V}$. In SALT, the overall amplitude of the light-curve at peak brightness, $x_{0}^{s}$, has no direct correspondent in BayeSN, which infers the distance jointly with the other parameters. $x_{1}^{s}$, $c^{s}$, and $x_{0}^{s}$ are all determined for each supernova, whereas $M_{0}$ and $M_{1}$ are population-level SEDs that are determined during the SALT training process.
@@
The dust-modeling framework introduced in Mandel et al. (2017); Brout & Scolnic (2021); Popovic et al. (2023b) makes some changes to the SALT and Tripp formulations. The distribution of SN Ia colours is comprised of a combination of a intrinsic, dust-free distribution - $c_{\rm int}$ - and a reddening component due to dust ($E_{\rm dust}$). The $c$ component in Equation 5 is then
@@
2.3 Differences between BayeSN and SALT: Dust
@@
  • BayeSN assumes a constant, achromatic scatter $\delta M^{s}$ around the distance-redshift relation which follows a Gaussian distribution, in addition to a chromatic scatter term. SALT does not explicitly include a grey scatter term equivalent to $\delta M^{s}$, though conventional cosmology analyses with SALT will approximate the grey scatter as a Gaussian.
@@ 3 Data
We make use of the Dark Energy Survey ‘5-year’ data release (DES-SN, (Flaugher et al., 2015; Sánchez et al., 2024). The DES observing program ran for 5 years, observing likely SNe Ia from $0.1<z<1.13$ with grizgriz bands. The host galaxy information for the DES supernovae is taken from co-added images (Wiseman et al., 2020), with environmental properties such as stellar mass and colour fit from the spectral energy distributions by Smith et al. (2020); Kelsey et al. (2023). A spectroscopic follow-up program by the Australian DES survey, OzDES (Childress et al., 2017; Lidman et al., 2020), provided spectroscopic redshifts of the host galaxies, which was paired with the state-of-the-art photometric classification program SuperNNova from Möller & de Boissière (2020). The resulting DES-SN sample is the largest single-telescope sample to-date, comprised of 1500 likely SNe Ia. A more comprehensive review of the SNe Ia used in the DES 5-year analysis can be found in Vincenzi et al. (2024); Möller et al. (2024).
@@ 3.1 Establishing the BayeDESN Sample
In short, these cuts give us the 393 SNe that define the BayeDESN sample. We additionally consider an alternative sample, the ‘Cosmo’ sample, which follows these requirements but additionally includes only SNe in the final DES5YR cosmology sample. The specific cuts applied to these samples, and the number of SNe they respectively cut, are detailed in Table 1.
@@ 4 Methodology
The focus of this work is to apply BayeSN to infer dust properties on SALT-based simulations, given the history of differing results concerning dust between the two. This will test BayeSN’s ability to accurately infer dust properties and potentially illustrate the reason for previous differences in results. We perform a suite of tests with simulated data, as detailed in sub-sections 4.1 and 4.2. We present the results of these tests in Section 5 then, in Section 6, we apply BayeSN to the real BayeDESN sample, to infer differences in dust and intrinsic properties on either side of the mass step.
@@ 4.1 Recovery of Simulated Parameters
Following Section 2.2, we simulate 50 simulacra of the DES-5YR dataset, varying supernova and dust properties according to Table 2. These simulations are then fit using the BayeSN model for hierarchical inference. We infer all global parameters including $W_{0},W_{1},\Sigma_{\epsilon}$ as well as the mean and standard deviation of the $R_{V}$ population, the population mean extinction $\tau_{A}$, and an achromatic intrinsic scatter $\sigma_{0}$; this allows the BayeSN model the flexibility to capture the variation in the SALT SED model while simultaneously inferring dust properties. In the cases where our simulated dust properties vary with host galaxy mass (see Table 2), we similarly split our hyperparameters to match the simulated ‘step’ at $10^{10}M_{*}$, where our dust distributions change. In these cases, we apply the ‘intrinsic SED difference’ model discussed in Section 2.1.1 to apply the most general model possible, with the exception of 5.2.1 where we attempt to constrain a simulated intrinsic mass step and therefore apply the ‘intrinsic mag difference’ model to directly infer a mass step parameter.
Systematic Test $R_{V}$ E(BV)E(B-V) $\beta_{\rm SN}$
Single $R_{V}$, E(BV)E(B-V) $\mathcal{N}(3,0.5)$ $\exp(0.080)$ 1
Double $R_{V}$, Single E(BV)E(B-V) $\mathcal{N}(2,0.5),\mathcal{N}(3,0.5)$ $\exp(0.080)$ 1
Double $R_{V}$, E(BV)E(B-V) $\mathcal{N}(2,0.5),\mathcal{N}(3,0.5)$ $\exp(0.125)$, $\exp(0.080)$ 1